OpenStudy (lgbasallote):

i need a demo of the exact DE method thingy... \[2xydx + (y^2 + x^2)dy = 0\]

OpenStudy (lgbasallote):

@Zarkon can you help me here if possible?

OpenStudy (amistre64):

integrate one of the "terms"

OpenStudy (lgbasallote):

doesnt it have that partial derivative thingy??

OpenStudy (amistre64):

yes

OpenStudy (lgbasallote):

could you teach me how to use that here?

OpenStudy (amistre64):

yes, which is your favorite side, right or left?

OpenStudy (lgbasallote):

uhh left

OpenStudy (amistre64):

then lets use the left: integrate 2xy dx

OpenStudy (lgbasallote):

with respect to which variable?

OpenStudy (lgbasallote):

oh x *facepalm*

OpenStudy (amistre64):

the dx says to do it to x

OpenStudy (lgbasallote):

yx^2 right?

OpenStudy (amistre64):

almost try again

OpenStudy (amistre64):

err, you might be right lol

OpenStudy (lgbasallote):

lol =)) so what's next?

OpenStudy (amistre64):

2y pulls out as a constant, and x goes to x^2/2 .. yep your good

OpenStudy (amistre64):

dont forget your +C tho

OpenStudy (lgbasallote):

i thought C can be put anywhere? o.O i guess this is a different approach...

OpenStudy (lgbasallote):

so it's yx^2 + c

OpenStudy (amistre64):

think of the +C as a function of y that gets treated as a constant with respect to x; say + g(y) instead

OpenStudy (amistre64):

yx^2 + g(y) ; now take the derivative with respect to y

OpenStudy (lgbasallote):

2xydx?

OpenStudy (amistre64):

yx^2 + g(y) wrt y doesnt go back down to "wrt x"

OpenStudy (lgbasallote):

oh sorry

OpenStudy (lgbasallote):

x^2 + 1

OpenStudy (lgbasallote):

oh no...x^2 only

OpenStudy (amistre64):

the derivative of g(y) wrt y = g'(y) right?

OpenStudy (lgbasallote):

oh wait...im confused @_@

OpenStudy (lgbasallote):

oh

OpenStudy (lgbasallote):

yeah...makes sense lol...so x^2 + g"(y)?

OpenStudy (lgbasallote):

^that'sone apostrophe only

OpenStudy (amistre64):

yx^2 + g(y) wrt y downs to: x^2 + g'(y) and this should equate to the right side that we left down there

OpenStudy (lgbasallote):

does it?

OpenStudy (amistre64):

x^2 + g'(y) = x^2 + y^2 g'(y) = y^2 ; integrate back up to get to g(y)

OpenStudy (lgbasallote):

im confused o.O

OpenStudy (lgbasallote):

why do we want g(y)?

OpenStudy (amistre64):

because g(y) is an unknown to us that needs to be determined to find a solution

OpenStudy (lgbasallote):

i see...

OpenStudy (lgbasallote):

so i'll integrate y^2?

OpenStudy (amistre64):

yep

OpenStudy (lgbasallote):

y^3 / 3

OpenStudy (amistre64):

+C which this time can just be a +C

OpenStudy (lgbasallote):

not g(x) anymore?

OpenStudy (amistre64):

we already have a f(x), its what we worked off of to begin with

OpenStudy (lgbasallote):

so...we have g(y) and f(x)...what now?

OpenStudy (amistre64):

\[yx^2 + \frac{1}{3}y^3+c\]is our results

OpenStudy (amistre64):

the idea is, up one side to find a function that can be down to the other side and compared to

OpenStudy (amistre64):

spose we started withe the right side: integrate x^2 + y^2 dy

OpenStudy (lgbasallote):

y^3/3 + x^2 y

OpenStudy (amistre64):

+c, which we regard as +f(x)

OpenStudy (amistre64):

derive that wrt x and compare it

OpenStudy (lgbasallote):

2xydx

OpenStudy (amistre64):

2xy + f'(x) = 2xy f'(x) = 0 ; integrate it back up to find f(x) f(x) = c

OpenStudy (amistre64):

either way we go we end up with:\[yx^2 + \frac{1}{3}y^3+c\]

OpenStudy (lgbasallote):

ahh i see..good method

OpenStudy (lgbasallote):

the method our teacher proposed was integrate both and then record all unique terms...seems the same will come out

OpenStudy (amistre64):

there are a few methods, im prone to this one tho :)